OpenStudy (anonymous):

if the area of a square is multiplied by 9, the area becomes 16 square inches. Find the length x of a side of the square.

OpenStudy (anonymous):

\[9x^2=16 \\x^2=\frac{16}{9} \\x=\frac43~~\text{inches} \]